what is the resultant of 12N force and 9 N force pulling an object and forming an angle of 60 degree with each other

Grade:12

2 Answers

Preetish Bhakne
4 years ago
Force is a vector quantity so we can add them vectorally i.e, 
Here f1=12 N
F2= 9 N
Angle= 60 degrees
Resultant = (f1+f2+2f1f2cos angle)^1/2
                 = (12+9+2(108)1/2)^1/2
                 = (129)^1/2
Arun
4 years ago
 
Dear student
 
As we know
 
R^2 = 12^2 + 9^2 + 2 * 12 * 9 * cos 60
 
R^2 = 144 + 81 + 108 =333
 
R = sqrt(333) = 18.24 Newton
 
Regards